Let f
f
be the function given by f(x)=∣∣x2−3∣∣⋅(x+0.5)(x2−3)(x+0.5)
f
(
x
)
=
|
x
2

3
|

(
x
+
0.5
)
(
x
2

3
)
(
x
+
0.5
)
. On which of the following open intervals is f
f
continuous?
1
Expert's answer
2019-09-27T09:46:43-0400

The function "|x^2-3|" is continuous on "\mathbb R", because it is composition of continuous on "\mathbb R" functions "x^2-3" and "|x|" (that is "|x^2-3|=g(h(x))" , where "h(x)=x^2-3" and "g(x)=|x|").

So "|x^2-3|(x+0.5)(x^2-3)(x+0.5)" is a continuous on "\mathbb R" function, because it is the product of the 4 continuous on "\mathbb R" functions.

Since "|x^2-3|(x+0.5)(x^2-3)(x+0.5)" is a continuous function on "\mathbb R", it is continuous on all subsets of "\mathbb R".